40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  TCS elog  Not logged in ELOG logo
Message ID: 251     Entry time: Wed May 11 19:46:21 2022
Author: Aidan, Radhika, Jordan 
Type: Lab Infrastructure 
Category: Vacuum chamber 
Subject: Leak cleaning on IR labs vacuum chamber - suspects 

[Aidan, Jordan, Radhika]

Radhika and Jordan identified some particulates (hair and flecks of foil) on the O-ring on the IR labs dewer. Additionally, we saw a scratch in the O-ring groove and a nick on the metal of the base of the dewer where it meets the O-ring. All were in the leaky vicinity previously identified by the He testing.

We set up a cradle to hold the dewer while we are working on it. Still needs vertical supports.

R & J replaced the O-ring with a new one with Crytox applied. 

Attachment 1: IMG_8812.jpg  555 kB  Uploaded Wed May 11 20:49:18 2022  | Hide | Hide all
IMG_8812.jpg
Attachment 2: IMG_8799.jpg  499 kB  Uploaded Wed May 11 20:49:31 2022  | Hide | Hide all
IMG_8799.jpg
Attachment 3: IMG_8789.jpg  633 kB  Uploaded Wed May 11 20:49:41 2022  | Hide | Hide all
IMG_8789.jpg
Attachment 4: IMG_8805.jpg  747 kB  Uploaded Wed May 11 20:50:55 2022  | Hide | Hide all
IMG_8805.jpg
ELOG V3.1.3-